OpenStudy (anonymous):

graphing inequalities: \[x+2y \le16\] graph this

OpenStudy (anonymous):

how would I graph this?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[0=-\frac{ 1 }{ 2 }x+8\]

OpenStudy (anonymous):

isn't slope intercept y=mx+b?

OpenStudy (praxer):

Let try it in another way.Okay ?

OpenStudy (anonymous):

k

OpenStudy (praxer):

let, x+2y=16 now dividing both sides by 16 we get, x/16+y/8=1 so, when x = 0, y=8 and y=0, x=16.  now for the inequality lets take that x=y=0 =>0 le 16 So, (0,0) can be a solution of the equation .

OpenStudy (praxer):